OpenStudy (anonymous):

HELP please :) A patient requires an IV of 0.9% saline solution, also known as normal saline solution. How much distilled water, to the nearest milliliter, must be added to 100 milliliters of a 6 % saline solution to produce normal saline?

OpenStudy (anonymous):

:)

OpenStudy (madhu.mukherjee.946):

:)

OpenStudy (anonymous):

:(

OpenStudy (madhu.mukherjee.946):

@Rushwr @ganeshie8

OpenStudy (madhu.mukherjee.946):

wait @TRICIACOLLIN i'm trying to help u

OpenStudy (anonymous):

Thanks... :))) go on. i thought i was snobbed. hihi

OpenStudy (madhu.mukherjee.946):

@Abhisar @aaronq

OpenStudy (madhu.mukherjee.946):

@chmvijay

OpenStudy (abhisar):

x % of a saline solution means that x grams of salts are dissolved in 100 ml of water.

OpenStudy (madhu.mukherjee.946):

okay but pls solve this problem

OpenStudy (abhisar):

Do you know the answer?

OpenStudy (anonymous):

sadly no :(

OpenStudy (abhisar):

Ok, no probs...just wanted to check.

OpenStudy (abhisar):

We can use, Initial volume \(\times\) Initial concentration = Final volume \(\times\) Final concentration

OpenStudy (madhu.mukherjee.946):

then

OpenStudy (abhisar):

Consider you need to add x ml of water then you can write \(\sf 0.06\times 100=0.009 \times (100+x)\) Solve for x

OpenStudy (abhisar):

I'll suggest you to cross check it cz my physical chem is rusted ...

OpenStudy (madhu.mukherjee.946):

is the answer 566.67 ml

OpenStudy (anonymous):

yes.. thank you! :)))) to both of you :)))

OpenStudy (anonymous):

A quadratic equation with real coefficients and leading coefficient 1, has x = -bi as a root. Write the equation in general form. :))